OpenStudy (geekfromthefutur):

I give medal and fan if you help me plz Solve 2x2 - 9x + 7 = 0 using an appropriate method. Show the steps of your work and explain why you chose the method used.

OpenStudy (geekfromthefutur):

@iambatman

OpenStudy (anonymous):

i have no clue sorry

OpenStudy (anonymous):

Factor/ quadratic equation

OpenStudy (anonymous):

Okay so this is a quadratic equation. \[2x ^{2}-9x+7=0\]

OpenStudy (geekfromthefutur):

Yea it is

OpenStudy (anonymous):

You can use this quadratic formula or you can factor normally, actually you can use complete the square as well, but I don't really suggest it for this problem. Quadratic formula: \[ x = \frac{ -b \pm \sqrt{b^2-4ac} }{ 2a }\]

OpenStudy (anonymous):



OpenStudy (geekfromthefutur):

Ok

OpenStudy (anonymous):

If you decide to factor you have to find a way to have the last two numbers in the parenthesis to multiply to 7 and add up to -9 after factoring out

OpenStudy (geekfromthefutur):

Ok

OpenStudy (geekfromthefutur):

@ganeshie8 im soo confused still

OpenStudy (anonymous):

I think it would be more ideal to factor this problem. Just plug and chug. What multiples to 7 since it's prime it's gonna be just 7 & 1 Try plugging them in until you get the same equation (2x(+or-)#)(x(+or-)#)

OpenStudy (geekfromthefutur):

Theres 2 answers to this equation ik that